OpenStudy (anonymous):

Each day at a manufacturing plant, a lot of 8 items are manufactured and three are known to be defective. A sample of 2 of these items is selected from the lot without replacement. What is the probability distribution for the number of defective parts? Also, describe the sample space by letting c indicate a non-defective and d indicate a defective part.

OpenStudy (anonymous):

@Zarkon

OpenStudy (anonymous):

@radar

OpenStudy (anonymous):

anyone, knows this problem

OpenStudy (kropot72):

\[P(0\ defectives)=\frac{3C0\times5C2}{8C2}=you\ can\ calculate\] \[P(1\ defective)=\frac{3C1\times5C1}{8C2}=you\ can\ caculate\] \[P(2\ defectives)=\frac{3C2\times5C0}{8C2}=you\ can\ calculate\]